OpenStudy (anonymous):

A cylinder has a volume of 175 cubic units and a height of 7 units. The diameter of the cylinder is

OpenStudy (goformit100):



OpenStudy (anonymous):

answer choices are 5units 10units 12.5units

OpenStudy (anonymous):

\[V=\pi r ^{2}h\] diameter=2r

OpenStudy (anonymous):

i dont understand what is the radius ?

OpenStudy (anonymous):

\[175=\pi r ^{2}h\] \[\pi=\frac{ 22 }{ 7 },h=7,find r\]

OpenStudy (anonymous):

ok so i got 21.98 when i multiplied 3.14 * 7 now what do i do

OpenStudy (anonymous):

\[\frac{ 22 }{ 7 }r ^{2}*7=175\] \[r ^{2}=\frac{ 175 }{22 },r=\sqrt{\frac{ 175 }{22 }}\] calculate r then diameter =2*r

OpenStudy (anonymous):

r=2.82 diameter=2*2.82=5.6 nearly